# Mathematical Language Models: A Survey

### v3

@@ Abstract.
In recent years, there has been remarkable progress in leveraging Language Models (LMs), encompassing Pre-trained Language Models (PLMs) and Large-scale Language Models (LLMs), within the domain of mathematics. This paper conducts a comprehensive survey of mathematical LMs, systematically categorizing pivotal research endeavors from two distinct perspectives: tasks and methodologies. The landscape reveals a large number of proposed mathematical LLMs, which are further delineated into instruction learning, tool-based methods, fundamental CoT techniques, and advanced CoT methodologies. In addition, our survey entails the compilation of over 60 mathematical datasets, including training datasets, benchmark datasets, and augmented datasets. Addressing the primary challenges and delineating future trajectories within the field of mathematical LMs, this survey is positioned as a valuable resource, poised to facilitate and inspire future innovation among researchers invested in advancing this domain.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Natural language generationccs: Computing methodologies Scene understandingccs: Computing methodologies Cognitive roboticsccs: Computing methodologies Cognitive scienceccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Mathematics stands as a foundational skill integral to human intelligence, wielding significance across diverse fields such as natural sciences, engineering, medicine, finance, computer science, and social sciences. Within the field of natural language processing (NLP), the development of computer models aimed at autonomously resolving mathematical word problems has captivated researchers since as early as 1963 (Feigenbaum et al., 1963; Bobrow et al., 1964; Briars and Larkin, 1984; Fletcher, 1985). We believe that addressing this problem is a potential way toward illuminating pathways for general reasoning mechanisms, consequently advancing the pursuit of general artificial intelligence (AI).
@@
The recent advent of LLMs, exemplified by OpenAI’s GPT-4 (OpenAI, 2023), has catalyzed an unforeseen surge in innovation, underscoring the multifaceted potential of AI within the domain of mathematics. These models ((Touvron et al., 2023a, b)) have demonstrated remarkable success across diverse Natural Language Processing (NLP) tasks, leveraging in-context learning (Min et al., 2022; Brown et al., 2020; Chen et al., 2022b) and instruction learning (Yang et al., 2023; Liu and Low, 2023). Recent studies by Wang et al. (Wang et al., 2023b) indicate that LLMs featuring over 100 billion parameters (e.g., GPT-3 (Brown et al., 2020) with 175 billion, PaLM (Chowdhery et al., 2022) with 540 billion) exhibit the capability to address intricate tasks by employing a chain-of-thought (CoT) mechanism (Wei et al., 2022) when furnished with a limited set of reasoning examples as demonstrations. Advancements in CoT frameworks (Chen et al., 2022a; Lu et al., 2023b; Zhang et al., 2023c; Fu et al., 2023) have been tailored to enhance mathematical performance, incorporating tools and programs (Gao et al., 2023; Drori et al., 2022; He-Yueya et al., 2023). An exemplar in this domain is LLEMMA (Azerbayev et al., [n. d.]), an openly accessible LLM specifically designed for mathematical tasks. LLEMMA showcases the capacity to generate self-contained textual solutions for mathematical problems and formulate formal proofs.
@@
We mainly review the most related surveys about mathematical reasoning and LMs (Min et al., 2023). Lu et al. (Lu et al., 2023c) delineate the landscape of deep learning applications specifically pertaining to mathematical reasoning. Qiao et al. (Qiao et al., 2023) predominantly delve into reasoning mechanisms centered around language models, encompassing arithmetic, commonsense, logical, symbolic, and multi-modal reasoning. Chu et al. (Chu et al., 2023) offer an exploration of studies focused on chain-of-thought (CoT) reasoning. Additionally, Qiu el al. (Qiu et al., 2020) and Zhao et al. (Zhao et al., 2023b) contribute comprehensive reviews elucidating the realm of existing PLMs and LLMs, respectively.
Diverging from these perspectives, our work delves into the burgeoning domain of mathematical language models, systematically summarizing the diverse range of studies and innovations in this field. In light of the demonstrated proficiency of LMs in comprehending, generating, and manipulating mathematical expressions, a multitude of advanced methodologies have emerged within the field of mathematics leveraging LMs. By shedding light on the current landscape, tasks, datasets, challenges, and future prospects, we aim to inspire and inform those who aspire to harness the power of language models to revolutionize mathematics and its myriad applications.
@@
Specifically, our categorization of mathematical tasks (§2) contain two primary domains: mathematical calculation (§2.1), consisting of arithmetic representation and arithmetic calculation, and mathematical reasoning (§2.2), consisting of problem-solving and theorem proving. Furthermore, our taxonomy of existing algorithms segregates them into PLMs-based approaches (§3), including autoregression (§3.1) and non-autoregression (§3.2) LMs, and LLMs-based methodologies (§4), which include instruction learning (§4.1), tool-based strategies (§4.2), fundamental CoT techniques (§4.3), and advanced CoT methodologies (§4.4). Moreover, we provide a comprehensive compilation of over 60 mathematical datasets (§5), systematically categorized into training (§5.1), benchmark (§5.2), and augmented datasets (§5.3). This classification aims to facilitate research by delineating the utility and applicability of these datasets within distinct research contexts. Lastly, our survey meticulously addresses the primary challenges and explores the future directions of this domain (§6), including faithfulness, multi-modal, uncertainty, evaluation, creation, application, and data scarcity. In conclusion (§7), we believe that this comprehensive survey will serve as a valuable asset for researchers and practitioners alike, seeking to push the boundaries of innovation in the field of mathematical language models.
@@ 2. Mathematical Tasks
In this section, we summarize the existing mathematical tasks into mathematical calculation and mathematical reasoning (Figure 1).
@@
Reasoning  (§2.2) [ Math Word Problem
Solving [ MathPrompter  (Imani et al., 2023), MetaMath  (Yu et al., 2023b), WizardMath  (Luo et al., 2023), MathAttack  (Zhou et al., 2023c), LLEMMA  (Azerbayev et al., [n. d.]) , leaf, text width=44.6em ] ] [ Math Question
Answering [ GEOS  (Seo et al., 2015), DROP  (Dua et al., 2019), Mathematics  (Saxton et al., 2019), Lila  (Mishra et al., 2022a) , leaf, text width=44.6em ] ] [ Theorem
@@ 2.1. Mathematical Calculation
The advent of Language Models (LMs) has ushered in a new era of exploration into their computational capabilities, particularly in arithmetic. Initially, these models demonstrated basic computational abilities by representing numbers in textual formats. With the increasing capabilities of LMs, it has been observed that they can acquire arithmetic skills through methods such as fine-tuning, even without carefully crafted numeric representations.
@@ 2.1.1. Arithmetic Representation
In the early stages of the research, numerical values were either omitted or oversimplified, treated as ordinary text, or categorized as “unknown” (UNK). This approach, however, proved inadequate for tasks. For instance, BERT performs five times worse when the answer is a number rather than a textual span over the DROP question-answering benchmark (Dua et al., 2019).
@@ 2.1.2. Arithmetic Calculation
Traditionally, it was presumed that LLMs could not accurately perform complex arithmetic, especially multiplication involving more than eight digits. However, recent approaches challenge this assumption. Zhou et al. (Zhou et al., 2022) apply specialized prompt engineering to enhance addition capabilities but note limitations in multiplication beyond seven digits. Jelassi et al. (Jelassi et al., 2023) investigate length generalization in basic arithmetic tasks using techniques like relative position embeddings and training set priming. ScratchpadGPT (Nye et al., 2021) demonstrates the effectiveness of pre-generating a Chain of Thought (CoT) before producing an answer in 8-digit addition tasks. Goat (Liu and Low, 2023) utilizes supervised instruction for fine-tuning elementary arithmetic operations with large integers. MathGLM (Yang et al., 2023) excels in intricate arithmetic tasks by pre-training on datasets that decompose complex arithmetic expressions into simpler steps. Through this process, it gradually generates answers and learns the rules of computation.
@@ 2.2.1. Math Problem Solving
In the field of math problem solving in artificial intelligence, math problem solving refers to the process of using algorithms, computational models, and increasingly LLMs to understand, explain, and solve mathematical problems. This approach to solving problems spans all levels from basic arithmetic to advanced mathematics, including but not limited to algebra, geometry, statistics, and calculus. In this subsection, we condense it into Math Word Problems (MWPs) and Math Question Answering (MQA).
Math Word Problem Solving.
@@
Math Word Problems (MWPs) refer to problems that present mathematical concepts and calculations in the form of written descriptions. Such questions usually include one or more story situations describing a series of events or conditions from which the solver needs to extract relevant mathematical information and apply appropriate mathematical principles to solve the problem. Over the years, using various efficient and intelligent algorithms to solve MWPs has also been a research focus in the fields of computer science and artificial intelligence (Bobrow et al., 1964).
Recent research has highlighted the growing capabilities of LLMs in the field of mathematical word problem solving, emphasizing the trend toward more nuanced and sophisticated AI-driven mathematical analysis. MathPrompter (Imani et al., 2023) uses LLM called GPT3 DaVinci to solve MWPs with excellent results, demonstrating the potential of LLMs to not only explain but also generate complex mathematical reasoning, reflecting a human-like understanding of complex problem sets.
Yuan et al. (Yuan et al., 2023a) investigated the interaction between various factors such as pre-training loss, amount of supervised data, and augmented data to improve the mathematical inference performance of LLMs. They found that pre-training loss is a more reliable performance metric and empirically found a log-linear relation between data amount and model performance, but it is worth mentioning that this improvement diminishes with better pre-training models. At the same time, they proposed Rejection sampling Fine-Tuning (RFT) to enhance performance by using supervised models to generate and collect correct reasoning paths, thereby achieving better inference generalization and more significant improvements. Meanwhile, the proficiency with which LLMs perform basic arithmetic tasks is critical to establishing the baseline capabilities of these models (Yuan et al., 2023b). This assessment is integral to understanding the limitations and potential of LLMs in future work when faced with mathematical challenges that require precise and logical calculations.
@@
MetaMath (Yu et al., 2023b) further extends the practicality of the LLMs, proposing a paradigm whereby the LLMs generate their mathematical problems, thereby creating a self-sustaining learning environment that encourages continuous improvement of the model’s problem-solving acuity. Similarly, WizardMath (Luo et al., 2023) explores new ways to enhance LLMs’ mathematical reasoning capabilities by reinforcing evolutionary instructions, indicating an important step towards autonomous self-improvement of artificial intelligence models. Instead, MathAttack (Zhou et al., 2023c) introduces a critical perspective by exploring the sensitivity of LLMs to specialized adversarial input designed to test their mathematical problem-solving abilities. Such investigations are crucial to developing more resilient and reliable LLMs that can withstand and adapt to a variety of problem-solving situations. LLEMMA (Azerbayev et al., [n. d.]), an open language model dedicated to mathematics, outlines ongoing efforts to make LLMs accessible and adaptable to a wide range of mathematical inquiries, from basic problem solving to advanced theorem proving.
The present findings confirm that these studies comprehensively depict the current research hotspots and future development potential directions of LLMs in the MWPs field. They emphasize that LLMs can have a transformative impact on the field of mathematics, providing innovative tools for educators, students, and researchers. As these models continue to develop, they are expected to revolutionize the way we approach and solve mathematical problems, ushering in a new era of mathematical reasoning and assisted education through artificial intelligence.
Math Question Answering.
@@
Math Question Answering (MQA) refers to the computational task where a system is required to understand and automatically solve math-related questions presented in natural language. These questions can range from simple arithmetic problems to complex high school or college-level mathematics, including algebra, calculus, and geometry. The challenge for MQA systems is to accurately interpret the text, convert it into an appropriate mathematical representation, perform the necessary computations, and generate a correct and often step-by-step solution, mimicking the problem-solving process of a human.
@@
Sachan et al. (Saxton et al., 2019) introduce a framework for evaluating neural architectures through a suite of math problems, highlighting the importance of algebraic generalization. This is complemented by GEOS (Seo et al., 2015), an innovative system that solves SAT geometry problems by integrating text understanding with diagram interpretation. The introduction of datasets and systems like Inter-GPS (Lu et al., 2021a), IconQA (Lu et al., 2021b), and PGDP5K (Hao et al., 2022; Zhang et al., 2022, 2023b) further advances the field, offering new benchmarks for abstract diagram understanding and visual language reasoning. Additionally, Scharpf et al. (Scharpf et al., 2022) indicate that unsupervised methods for formula labeling contribute to the automated understanding of mathematical documents. Collectively, these works demonstrate the evolving landscape of MQA, where the synergy between natural language processing, computer vision, and symbolic reasoning opens new avenues for educational technology and AI research.

### v4

@@ Abstract.
In recent years, there has been remarkable progress in leveraging Language Models (LMs), encompassing Pre-trained Language Models (PLMs) and Large-scale Language Models (LLMs), within the domain of mathematics. This paper conducts a comprehensive survey of mathematical LMs, systematically categorizing pivotal research endeavors from two distinct perspectives: tasks and methodologies. The landscape reveals a large number of proposed mathematical LLMs, which are further delineated into instruction learning, tool-based methods, fundamental CoT techniques, advanced CoT methodologies and multi-modal methods. To comprehend the benefits of mathematical LMs more thoroughly, we carry out an in-depth contrast of their characteristics and performance. In addition, our survey entails the compilation of over 60 mathematical datasets, including training datasets, benchmark datasets, and augmented datasets. Addressing the primary challenges and delineating future trajectories within the field of mathematical LMs, this survey is poised to facilitate and inspire future innovation among researchers invested in advancing this domain.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Natural language generationccs: Computing methodologies Scene understandingccs: Computing methodologies Cognitive roboticsccs: Computing methodologies Cognitive scienceccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Mathematics stands as a foundational skill integral to human intelligence, wielding significance across diverse fields such as natural sciences, engineering, medicine, finance, computer science, and social sciences. Within the field of natural language processing (NLP), the development of computer models aimed at autonomously resolving mathematical word problems has captivated researchers since as early as 1963 (Feigenbaum et al., 1963; Bobrow et al., 1964; Briars and Larkin, 1984; Fletcher, 1985). We believe that addressing this problem is a potential way toward illuminating pathways for general reasoning mechanisms, consequently advancing the pursuit of artificial general intelligence (AGI).
@@
The recent advent of LLMs, exemplified by OpenAI’s GPT-4 (OpenAI, 2023), has catalyzed an unforeseen surge in innovation, underscoring the multifaceted potential of AI within the domain of mathematics. These models ((Touvron et al., 2023a, b)) have demonstrated remarkable success across diverse Natural Language Processing (NLP) tasks, leveraging in-context learning (Min et al., 2022; Brown et al., 2020; Chen et al., 2022b) and instruction learning (Yang et al., 2023; Liu and Low, 2023). Recent studies by Wang et al. (Wang et al., 2023d) indicate that LLMs featuring over 100 billion parameters (e.g., GPT-3 (Brown et al., 2020) with 175 billion, PaLM (Chowdhery et al., 2022) with 540 billion) exhibit the capability to address intricate tasks by employing a chain-of-thought (CoT) mechanism (Wei et al., 2022) when furnished with a limited set of reasoning examples as demonstrations. Advancements in CoT frameworks (Chen et al., 2022a; Lu et al., 2023c; Zhang et al., 2023e; Fu et al., 2023) have been tailored to enhance mathematical performance, incorporating tools and programs (Gao et al., 2023; Drori et al., 2022; He-Yueya et al., 2023). Moreover, openly accessible LLMs are specifically designed for mathematical tasks, such as LLEMMA (Azerbayev et al., [n. d.]), Qwen-Math (Yang et al., 2024) and InternLM-Math (Ying et al., 2024). Recently, o1 (OpenAI, 2024) and o3 obtained state-of-the-art performance on mathematical reasoning by integrating reinforcement learning with the Monte Carlo tree.
@@
We mainly review the most relevant surveys about mathematical reasoning and LMs (Min et al., 2023). For surveys focused on mathematical reasoning, Lu et al. (Lu et al., 2023d) delineate the landscape of deep learning applications specifically pertaining to mathematical reasoning (DL4Math). Qiao et al. (Qiao et al., 2023) predominantly delve into general reasoning mechanisms centered around language models, encompassing arithmetic, commonsense, logical, symbolic, and multi-modal reasoning (LM4Reasoning). Additionally, Chu et al. (Chu et al., 2023) offer an exploration of studies focused on CoT reasoning. For surveys related to LMs, Qiu et al. (Qiu et al., 2020) and Zhao et al. (Zhao et al., 2023b) contribute comprehensive reviews elucidating the realm of existing PLMs and LLMs, respectively. Yan et al. (Yan et al., 2024) summary the related studies about multimodal LLMs-based mathematical reasoning (MMLM4Math).
In contrast to these existing works, our study shifts focus to the emerging field of mathematical language models (MLMs), presenting a systematic and in-depth examination of the novel methods, tasks, and datasets that have emerged specifically within this domain. While previous surveys primarily address general reasoning capabilities of LMs or broader trends in LMs, our work distinguishes itself by offering a detailed exploration of how LMs are being uniquely applied to advanced mathematical tasks. We not only discuss the successes and challenges of existing approaches but also identify critical gaps and emerging trends that are shaping the future of mathematical reasoning. By providing a comprehensive overview of mathematical language models, including their applications, challenges, and future directions, our work aims to inspire further research and innovation in this rapidly developing area. We believe that highlighting the intersections between mathematics and LMs will encourage new approaches for solving complex mathematical problems, and ultimately, contribute to the broader goal of leveraging language models to revolutionize mathematics and its real-world applications.
@@
Specifically, our categorization of mathematical tasks (§2) contain two primary domains: mathematical calculation (§2.1), consisting of arithmetic representation and arithmetic calculation, and mathematical reasoning (§2.2), consisting of math problem-solving and theorem proving. Furthermore, our taxonomy of existing algorithms segregates them into PLMs-based approaches (§3), including autoregression (§3.1) and non-autoregression (§3.2) LMs, and LLMs-based methodologies (§4), which include instruction learning (§4.1), tool-based strategies (§4.2), fundamental CoT techniques (§4.3), advanced CoT methodologies (§4.4) and multi-modal methods (§4.5). Moreover, we provide a comprehensive compilation of over 60 mathematical datasets (§5), systematically categorized into training (§5.1), benchmark (§5.2), and augmented datasets (§5.3). This classification aims to facilitate research by delineating the utility and applicability of these datasets within distinct research contexts. Also, we analyze and discuss the characteristics and performance of various mathematical LMs (§6). Lastly, our survey meticulously addresses the primary challenges and explores the future directions of this domain (§7), including faithfulness, multi-modal, uncertainty, evaluation, creation, application, and data scarcity. In conclusion (§8), we believe that this comprehensive survey will serve as a valuable asset for researchers and practitioners alike, seeking to push the boundaries of innovation in the field of mathematical language models.
@@ 2. Mathematical Tasks
In this section, we summarize the existing mathematical tasks into mathematical calculation and mathematical reasoning (Figure 1, 2).
@@
Reasoning  (§2.2) [ Math Problem
Solving [ MathPrompter  (Imani et al., 2023), MetaMath  (Yu et al., 2023b), WizardMath  (Luo et al., 2023), MathAttack  (Zhou et al., 2023c), LLEMMA  (Azerbayev et al., [n. d.]), GEOS  (Seo et al., 2015),
DROP  (Dua et al., 2019), Mathematics  (Saxton et al., 2019), Lila  (Mishra et al., 2022a)
, leaf, text width=44.6em ] ] [ Theorem
@@
Figure 2. Examples of mathematical tasks, where (a) illustrates responses with different representation types.
@@ 2.1. Mathematical Calculation
The advent of LMs has ushered in a new era of exploration into their computational capabilities, particularly in arithmetic. Initially, these models demonstrated basic computational abilities by representing numbers in textual formats. With the increasing capabilities of LMs, it has been observed that they can acquire arithmetic skills through methods such as fine-tuning, even without carefully crafted numeric representations.
@@ 2.1.1. Arithmetic Representation
In the early stages of the research, numerical values were either omitted or oversimplified, treated as ordinary text, or categorized as “unknown”. These approaches, however, proved inadequate for mathematical tasks. For instance, BERT performs five times worse when the answer is a number rather than a textual span on the DROP benchmark (Dua et al., 2019).
@@ 2.1.2. Arithmetic Calculation
Traditionally, it was presumed that LLMs could not accurately perform complex arithmetic, especially multiplication involving more than eight digits. However, recent approaches challenge this assumption. Zhou et al. (Zhou et al., 2022) apply specialized prompt engineering to enhance addition capabilities but note limitations in multiplication beyond seven digits. Jelassi et al. (Jelassi et al., 2023) investigate length generalization in basic arithmetic tasks using techniques like relative position embeddings and training set priming. ScratchpadGPT (Nye et al., 2021) demonstrates the effectiveness of pre-generating a Chain of Thought (CoT) before producing an answer in 8-digit addition tasks. Goat (Liu and Low, 2023) utilizes supervised instruction for fine-tuning elementary arithmetic operations with large integers. MathGLM (Yang et al., 2023) excels in intricate arithmetic tasks by pre-training on datasets that decompose complex arithmetic expressions into simpler steps.
@@ 2.2.1. Math Problem Solving
Math problem solving refers to the process of using algorithms, computational models, and increasingly LLMs to understand, explain, and solve mathematical problems. These approaches to solving problems spans all levels from basic arithmetic to advanced mathematics, including but not limited to algebra, geometry, statistics, and calculus. Generally, mathematical problem solving can be categorized into two primary forms: the first involves contextualized problems, where the solver must comprehend the given scenario, translate it into a purely mathematical problem, and then proceed to solve it with solutions; the second form presents the solver with a direct mathematical problem, requiring only the provision of a specific answer.
@@
Math Word Problems (MWPs) refer to problems that present mathematical concepts and calculations in the form of written descriptions. Such questions usually include one or more story situations that describe a series of events or conditions from which the solver needs to extract relevant mathematical information and apply appropriate mathematical principles to solve the problem. Over the years, the use of various efficient and intelligent algorithms to solve MWPs has also been a research focus in the fields of computer science and artificial intelligence (Bobrow et al., 1964). Recent research has highlighted the growing capabilities of LLMs in MWPs task, emphasizing the trend toward more nuanced and sophisticated AI-driven mathematical analysis. MathPrompter (Imani et al., 2023) uses a GPT3 DaVinci to solve MWPs with excellent results, demonstrating the potential of LLMs to not only explain but also generate complex mathematical reasoning. Yuan et al. (Yuan et al., 2023a) investigated the interaction between various factors such as pre-training loss, amount of supervised data, and augmented data to improve the mathematical inference performance of LLMs. They proposed Rejection sampling Fine-Tuning (RFT) to enhance performance by using supervised models to generate and collect correct reasoning paths, thereby achieving better inference generalization.
@@
MetaMath (Yu et al., 2023b) further extends the practicality of LLMs, proposing a paradigm by which LLMs generate their mathematical problems, thus creating a self-sustaining learning environment that encourages continuous improvement of the model’s problem-solving acuity. Similarly, WizardMath (Luo et al., 2023) explores new ways to enhance LLMs’ mathematical reasoning capabilities by reinforcing evolutionary instructions, indicating an important step towards the autonomous self-improvement of models. Instead, MathAttack (Zhou et al., 2023c) introduces a critical perspective by exploring the sensitivity of LLMs to specialized adversarial input designed to test their mathematical problem-solving abilities. Such investigations are crucial to developing more resilient and reliable LLMs that can withstand and adapt to a variety of problem-solving situations. Furthermore, LLEMMA (Azerbayev et al., [n. d.]), an open-source LM dedicated to mathematics, outlines ongoing efforts to make LLMs accessible to a wide range of mathematical inquiries, from basic problem solving to advanced theorem proving.
@@
Correspondingly, if the problems are non-contextualized and purely mathematical, the focus shifts more towards the Math Question-Answer (MQA) system’s ability to understand and automatically solve the related computational tasks. These questions can range from simple arithmetic problems to complex high-school or college-level mathematics, including algebra, calculus, and geometry. The challenge for MQA systems is to accurately interpret the text, convert it into an appropriate mathematical representation, perform the necessary computations, and generate a correct and often step-by-step solution, mimicking the problem-solving process of a human.
@@
Sachan et al. (Saxton et al., 2019) introduce a framework for evaluating neural architectures through a suite of math problems, highlighting the importance of algebraic generalization. This is complemented by GEOS (Seo et al., 2015), an innovative system that solves SAT geometry problems by integrating text understanding with diagram interpretation. The introduction of datasets and systems like Inter-GPS (Lu et al., 2021a), IconQA (Lu et al., 2021b), and PGDP5K (Hao et al., 2022; Zhang et al., 2022, 2023d) further advances the field, offering new benchmarks for abstract diagram understanding and visual language reasoning. Additionally, Scharpf et al. (Scharpf et al., 2022) indicate that unsupervised methods for formula labeling contribute to the automated understanding of mathematical documents. Collectively, these works demonstrate the evolving landscape of MQA, where the synergy between natural language processing, computer vision, and symbolic reasoning opens new avenues for educational technology and AI research.